To expand the camera business, you decide to advertise your new sports camera model named FlyPro. After some research, it appears two options are available to you:

Billboards: you can rent 5 of those at $20,000 each per month, and all together roughly 100,000 people in that month will see your ads on billboards. Because the billboards are located in upscale areas of the town, about 80% of those who see the ad will buy your FlyPro model of sports camera.

Google ad: you can place your ad on Google for $1,000,000 per month so that it will be visible to about 3 million individuals all around the world. However, because many of those individuals reside in countries where they can't purchase or access your camera, you estimate that only 1% of those who see your ad will actually buy the product.

Given the above scenario, which statement below is true?

(recall average cost of advertising is how much you spent on advertising per buyer, that is, per actual customer)

A) The per customer cost of reaching potential customers using billboards is $0.1

B) The average cost of advertising using Google is exactly $1

C) While the per customer cost of reaching potential customers is lower using billboards, the average cost of advertising with billboards is higher

D) While the per customer cost of reaching potential customers is lower using Google, the average cost of advertising with Google is higher
Answer #1
Average cost of advertising with billboards 1 =(20000*5)/100000
Per customer cost of reaching potential customers using billboards 1.25 =(20000*5)/(100000*80%)
Average cost of advertising with Google 0.33 =1000000/3000000
Per customer cost of reaching potential customers using Google 33.33 =1000000/(3000000*1%)
While the per customer cost of reaching potential customers is lower using billboards, the average
cost of advertising with billboards is higher
Option C is correct